OpenStudy (anonymous):

|3x+10|-|2x+1|=4 solve this please

OpenStudy (anonymous):

one possibility \[(-(3x+10))-(-(2x+1))\]

OpenStudy (anonymous):

first possible solution was -13 second possiblity \[((3x+10))-((2x+1))==4\] =5 \[(-(3x+10))-((2x+1))==4\] -3 \[((3x+10))-(-(2x+1))==4\] -7/5

OpenStudy (anonymous):

try -13, 5, -3, 7/5 only some will work

OpenStudy (anonymous):

5 and -3 will not work

myininaya (myininaya):

4 cases to look at: ------------------------------------------------- 3x+10>0 => |3x+10|=3x+10 (x>-10/3) 2x+1>0=> |2x+1|=2x+1 (x>-1/2) ()******** ----|--------|-------- -10/3 -1/2 ------------------------------------------------- 3x+10<0 => |3x+10|=-3x+10 (x<-10/3) 2x+1>0 => |2x+1|=2x+1 (x>-1/2) ****() ()********** ----|---------|--------- -10/3 -1/2 ------------------------------------------------- 3x+10>0 => |3x+10|=3x+10 (x>-10/3) 2x+1<0 => |2x+1|=-(2x+1) (x<-1/2) ()***********() ---|----------|-------- -10/3 -1/2 ------------------------------------------------- 3x+10<0 => |3x+10|=-(3x+10) (x<-10/3) 2x+1<0 => |2x+1|=-(2x+1) (x<-1/2) ***() ---|--------|------- -10/3 -1/2 -------------------------------------------------- 4 different cases: Case 1: (3x+10)-(2x+1)=4 3x+10-2x-1=4 x+9=4 x=-5 |3(-5)+10|-|2(-5)+1|=5-9=-4 also x=-5 is not in the range that we had above ------------------------------------- Case 2: -(3x+10)-(2x+1)=4 -3x-10-2x-1=4 -5x-11=4 -5x=15 x=-3 |3(-3)+10|-|2(-3)+1|= 1-5=-4 not 4 also not in range for case 2 ------------------------------------- Case 3: (3x+10)--(2x+1)=4 5x+11=4 5x=-7 x=-7/5 |3(-7/5)+10|-|2(-7/5)+1|=4 YAY we found a solution! ------------------------------------- Case 4: -(3x+10)+(2x+1)=4 -3x-10+2x+1=4 -x-9=4 -x=13 x=-13 |3(-13)+10|-|2(-13)+1|=29-25= 4 YAY we have another solution -------------------------------------- Sorry for delay. I was having an allergy attack.

OpenStudy (anonymous):

Hope you are feeling all right